OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point. y sin12x = x cos2y, (π/2, π/4)

OpenStudy (anonymous):

hint: product rule and chain rules would be needed.

OpenStudy (anonymous):

what do you get? I got 3pi(x-pi/2) -3pi ( y-pi/4). don't know it 's right or wrong

OpenStudy (anonymous):

\[ (y)(12\cos(12x))+\sin(12x){dy\over dx}=(x)(-2\sin(2y)){dy\over dx}+\cos(2y) \]